OpenStudy (anonymous):

Subtract. Simplify by collecting like radical terms if possible. 6sqrt48 -2sqrt3=

OpenStudy (anonymous):

\[6 \sqrt48 -2\sqrt3=\]

OpenStudy (anonymous):

22sqrt{3}

OpenStudy (anonymous):

what's the {}?

OpenStudy (anonymous):

so \[\sqrt[22]{3}\] or \[22\sqrt 3\] ?